A Bibliography of Publications of Ulrich Ruede

Ulrich Ruede Fakultaet fuer Mathematik

Technische Universitaet Chemnitz-Zwickau

D-09009 Chemnitz Germany

Tel: +49 - 0371 - 561 -2159 FAX: ?n/a?

E-mail: ruede@mathematik.tu-chemnitz.de (Internet)

15 March 2022 Version 0.12

Abstract

This bibliography records publications of Ulrich Ruede.

Title word cross-reference

-Solution [GHRS92a, GHRS92b].

abstraction [Rüd93a]. Accurate [Rüd88a, KR04]. Adaptive [MR89, Rüd93f, Rüd91a, Rüd92a, Rüd94, PR93, Rüd92c, Rüd92d, Rüd93d]. Algebraic [MRZ88, RR93]. Algorithmen [ZFR88]. algorithms [Rüd93a]. AMG [MRZ88, RR93]. Analysis [BR92]. Anwendung [Rüd85]. Application [MRZ88]. Aspekte [ZFR88]. Augmented [DLP^{+}21].

based [Rüd93b]. Basis [Rüd92b]. Behandlung [Rüd88b]. Berechnung [Rüd85, SR90]. Betriebssystem [ZFR88]. Betriebssystem- [ZFR88]. Block [DLP^{+}21]. Boundary [BGR92].

Cimmino [DLP^{+}21]. Combination [BGR92, GHRS92a, GHRS92b]. Composite [MR89]. Computation [Rüd88a]. Computational [JKR08, Rue02, Rüd93d]. computations [JR94]. computing [KR04]. Constrained [MRZ88]. constructing [Rüd93c]. Convergence [MR89]. Corners [Rüd89]. Corrections [Rüd89]. cycle [Rüd92d].

Data [Rüd92a, Rüd93a]. Development [FR87, JRS88]. Differential [MR90, PR93]. Differentialgleichungen [Rüd88b]. digitalen [Rüd85]. Discretizations [Rüd86].

Ebene [FR90]. Editorial [RR03]. Effect [Rüd89]. efficiency [Rüd92c]. Element [RZ92, JR94, Rüd93c]. Eliminating [Rüd89]. Elliptic [BGR92, MR90, Rüd91b, KR04]. elliptischen [Rüd88b]. Engineering [JKR08, Rue02, ZFR88]. Equation [Rüd88a]. Equations [MR90, Rüd91b, PR93]. Erlangen [Rue02]. Erlangen-Nuremberg [Rue02]. Error [Rüd93b]. estimators [Rüd93b]. Expansion [JRS88]. Extensions [DLP^{+}21]. Extrapolation [BRSZ94, BGR92, KR04, Rüd87, Rüd91b, Rüd92b, Rüd93c, Rüd93e, JR94].

Fast [MR89]. Finite [MR89, RZ92, JR94, Rüd93c]. Full [DLP^{+}21]. Fully [Rüd93f, Rüd92c, Rüd92d].

Gauß [PR93]. Grid [BRSZ94, BGR92, GHRS92a, GHRS92b, MR89, Rüd93e]. grids [PR93].

Hierarchical [Rüd92b]. Higher [MR90, Rüd91a, Rüd93c]. Höhenmodellen [Rüd85].

Implicit [JR94]. Intensive [BR92]. Issue [JKR08]. Iterative [Rüd92a, Rüd94].

klassischer [SR90]. Kommando [FR90]. Konzepte [FR90].

Laplace [Rüd88a]. Layout [RR93]. Local [MR90, Rüd89].

Machines [GHRS92a]. Macro [JRS88]. Mathematical [Rüd93d]. Matrix [SR90]. Matrix-Multiplikationen [SR90]. Mehrgittermethode [Rüd85]. Method [DLP^{+}21, MR89, RZ86a, Rüd92b, RZ92, Rüd94, Rüd92c, Rüd92d]. Methods [AR91, BRSZ94, MR90, RR93, Rüd93f, Rüd86, RZ86b, Rüd87, Rüd91a, Rüd92a, Rüd93e, JR94, Rüd93c, Rüd93d]. Multigrid [AR91, MRZ88, RR93, Rüd93f, Rüd86, RZ86a, RZ86b, Rüd87, Rüd91a, Rüd92c, Rüd92d]. Multilevel [Rüd92a, Rüd93e, Rüd94, JR94, PR93, Rüd93a, Rüd93d]. Multiple [Rüd87]. Multiplikationen [SR90]. Multiprocessor [GHRS92a].

Networks [GHRS92a, GHRS92b]. Numerical [FR87]. Numerically [BR92]. numerischen [Rüd88b]. Nuremberg [Rue02].

Operating [FR87]. Optimization [BR92, MRZ88, RR93]. Order [MR90, Rüd91a, Rüd93c].

Parabolic [BRSZ94]. Parallel [FR87, GHRS92a, GHRS92b]. parallelen [FR90, ZFR88]. Partial [MR90, PR93]. partiellen [Rüd88b]. PDEs [GHRS92a, GHRS92b]. Performance [BR92]. Photogrammetrie [Rüd85]. Poisson’s [Rüd88a]. Pollution [Rüd89]. Preconditioning [GHRS92a, GHRS92b]. Problems [BRSZ94, BGR92, KR04]. Programmierung [FR90]. Programs [BR92, Rue02].

Quadratic [MRZ88].

Rank [DLP^{+}21]. Rectangular [DLP^{+}21]. Reentrant [Rüd89]. Refinement [MR90]. Related [Rüd91b]. relaxation [PR93]. robustness [Rüd92c].

Schnellere [SR90]. Science [JKR08]. Scientific [JRS88]. Singular [Rüd88a, KR04]. Singularitäten [Rüd88b]. Singularities [RZ86a, RZ92]. Software [FR87, JRS88, ZFR88]. Software-Engineering-Aspekte [ZFR88]. Solution [DLP^{+}21, GHRS92a, GHRS92b, PR93]. Solutions [Rüd88a, KR04]. Solvers [Rüd92a]. Solving [Rüd91b]. Sparse [BRSZ94, BGR92, GHRS92a, GHRS92b, Rüd93e, PR93]. Sparse-Grid-Preconditioning [GHRS92a, GHRS92b]. Special [JKR08]. splittings [Rüd93b]. stable [Rüd93b]. Structures [Rüd92a]. Support [FR87]. System [FR87]. Systematic [JRS88]. Systems [DLP^{+}21].

tau [Rüd87]. tau-Extrapolation [Rüd87]. Technique [GHRS92a, GHRS92b]. Techniques [BGR92, Rüd91b, KR04, Rüd93a, Rüd93c, Rüd93d]. Theory [MR89]. Tool [JRS88]. Toolbox [AR91]. Treatment [RZ86a, RZ92].

University [Rue02]. Unix [FR90]. Unix-Kommando-Ebene [FR90].

V [Rüd92d]. V-cycle [Rüd92d]. Value [BGR92]. Volume [MR89].

Werkzeuge [FR90]. Workbench [RZ86b]. Workstation [GHRS92a, GHRS92b].

zur [FR90, Rüd85, Rüd88b].

References

Arbesmeier:1991:TMM

[AR91] M. Arbesmeier and U. Rüde. A toolbox for multigrid methods. Bericht I-9136, Institut für Informatik, TU München, Septem-

ber 1991. URL file://www.tu-
chemnitz.de/pub/Local/mathematik/
Ruede/mgwb.ps.Z.

Bungartz:1992:ECS

[BGR92] H. Bungartz, M. Griebel, and U. Rüde. Extrapolation, combination and sparse grid techniques for elliptic boundary value problems. SFB Bericht 342/10/92 A, Institut für Informatik, TU München, May 1992. to be published in Computer Methods in Applied Mechanics and Engineering (1994).

Bonk:1992:PAO

[BR92] T. Bonk and U. Rüde. Performance analysis and optimization of numerically intensive programs. SFB Bericht 342/26/92 A, Institut für Informatik, TU München, November 1992. URL

file://www.tu-chemnitz.de/
pub/Local/mathematik/Ruede/
performance.ps.Z.

Balder:1994:SGE

[BRSZ94] R. Balder, U. Rüde, S. Schneider, and C. Zenger. Sparse grid and extrapolation methods for parabolic problems. In Proceedings of the 10th International Conference on Computational Methods in Water Resources, Heidelberg, 19.- 22. Juli 1994, 1994. URL

file://www.tu-chemnitz.de/
pub/Local/mathematik/Ruede/
heidel94.ps.Z.

Dumitrasc:2021:EAB

[DLP^{+}21] Andrei Dumitrasc, Philippe Leleux, Constantin Popa, Ulrich Ruede, and Daniel Ruiz. Extensions of the augmented block Cimmino method to the solution of full rank rectangular systems. SIAM Journal on Scientific Computing, 43(5):S516– S539, ???? 2021. CODEN SJOCE3. ISSN 1064-8275 (print), 1095-7197 (electronic).

Foessmeier:1987:OSS

[FR87] R. Fößmeier and U. Rüde. Operating system support for parallel numerical software development. Bericht I-8712, Institut für Informatik, TU München, October 1987.

Foessmeier:1990:KWP

[FR90] R. Fößmeier and U. Rüde.

Konzepte und Werkzeuge zur parallelen Programmierung auf der Unix-Kommando-Ebene. unix/mail, 8(2):66–73, 1990.

Griebel:1992:CTPa

[GHRS92a] M. Griebel, W. Huber, U. Rüde, and T. Störtkuhl. The combination technique for parallel sparse-grid-preconditioning and - solution of PDEs on multiprocessor machines and workstation networks. SFB Bericht 342/11/92 A, Institut für Informatik, TU München, May 1992.

Griebel:1992:CTPb

[GHRS92b] M. Griebel, W. Huber, U. Rüde, and T. Störtkuhl. The combination technique for parallel sparse-grid-preconditioning or - solution of PDEs on workstation networks. In L. Bougé, M. Cosnard, Y. Robert, and D. Trystram, editors, Parallel Processing: CONPAR 92 – VAPP V, volume 634 of Lecture Notes in Computer Science, pages 217–228. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1992. Proceedings of the Second Joint International Conference on Vector and Parallel Processing, Lyon, France, September 1–4, 1992.

Johnson:2008:SIC

[JKR08] Chris Johnson, David Keyes, and Ulrich Ruede. Special issue on computational science and engineering. SIAM Journal on Scientific Computing, 30(6):vii, ???? 2008. CODEN SJOCE3. ISSN 1064-8275 (print), 1095-7197 (electronic).

Jung:1994:IEM

[JR94] M. Jung and U. Rüde. Implicit extrapolation methods for multilevel finite element computations. In T. Manteuffel, editor, Preliminary Proceedings of the Colorado Conference on Iterative Methods, Breckenridge, Colorado, April 4-10, 1994, 1994. URL

                                        file://www.tu-chemnitz.de/
                                        pub/Local/mathematik/Ruede/
                                        impl_extrapolation.ps.Z.

Jaensch:1988:MET

[JRS88] Christian R. Jaensch, Ulrich Ruede, and Klaus Schnepper. Macro expansion, a tool for the systematic development of scientific software. Report I-8814, Institut für Informatik, TU München, München, West Germany, November 1988.

Koestler:2004:ETC

[KR04] H. Koestler and U. Ruede. Extrapolation techniques for computing accurate solutions of elliptic problems with singular solutions. In Computational science— ICCS 2004. Part IV, volume 3039 of Lecture Notes in Comput. Sci., pages 410–417. Springer, Berlin, 2004.

McCormick:1989:FVC

[MR89] S. McCormick and U. Rüde. A finite volume convergence theory for the fast adaptive composite grid method. Technical report, University of Colorado at Denver, 1989. To be published in Applied Numerical Mathematics (14) 1994, Elsevier.

McCormick:1990:LRH

[MR90] S. McCormick and U. Rüde.

On local refinement higher order methods for elliptic partial differential equations. International Journal of High Speed Computing, 2(4):311–334, 1990. Also available as TU-Bericht I-9034.

Muszynski:1988:AAM

[MRZ88] P. Muszynski, U. Rüde, and C. Zenger. Application of algebraic multigrid (AMG) to constrained quadratic optimization. Bericht I-8801, Institut für Informatik, TU München, January 1988.

Pflaum:1993:GAR

[PR93] C. Pflaum and U. Rüde. Gauß’ adaptive relaxation for the multilevel solution of partial differential equations on sparse grids. SFB-Bericht 342/13/93 A, Institut für Informatik, TU München, September 1993. URL

                                        file://www.tu-chemnitz.de/
                                        pub/Local/mathematik/Ruede/

gauss.ps.Z. to appear in the proceedings of the 2nd Gauß Symposium, Munich, Aug. 2–7, 1993.

Regler:1993:LOA

[RR93] H. Regler and U. Rüde. Layout optimization with algebraic multigrid methods (AMG). In Proceedings of the Sixth Copper Mountain Conference on Multigrid Methods, Copper Mountain, April 4-9, 1993, Conference Publication. NASA, 1993. URL

                                        file://www.tu-chemnitz.de/
                                        pub/Local/mathematik/Ruede/
                                        amg.ps.Z.

Renaut:2003:E

[RR03] Rosemary A. Renaut and Ulrich Ruede. Editorial. Future Generation Computer Systems, 19 (8):1265, November 2003. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Ruede:1985:AMB

[Rüd85] U. Rüde. Anwendung der Mehrgittermethode zur Berechnung von digitalen Höhenmodellen in der Photogrammetrie. Bericht I-8525, Institut für Informatik, TU München, November 1985.

Ruede:1986:DMM

[Rüd86] U. Rüde. Discretizations for multigrid methods. In W. Hackbusch and U. Trottenberg, editors, Multigrid Methods: Special Topics and Applications, Papers presented at the 2nd European Conference on Multigrid Methods, October 1-4, 1985, volume 110 of GMD Studien, pages ??–?? (of 178). Cologne, May 1986. ISBN 3-88457-110-9. LCCN QA377.E87 1985. Also available as TU-Bericht I-8519.

Ruede:1987:MTE

[Rüd87]  U. Rüde.  Multiple tau-

extrapolation for multigrid methods. Bericht I-8701, Institut für Informatik, TU München, January 1987.

Ruede:1988:ACS

[Rüd88a] U. Rüde. On the accurate computation of singular solutions of Laplace’s and Poisson’s equation. In S. F. McCormick, editor, Multigrid Methods: Theory, Applications, Supercomputing: Proceedings of the Third Copper Mountain Conference on Multigrid Methods, April 5-10, 1987, pages ??–?? (of xiv + 644). Marcel Dekker, New York, NY, USA, 1988. ISBN 0-8247-7979-7. LCCN QA377 .M9431 1988.

Ruede:1988:NBS

[Rüd88b] U. Rüde. Zur numerischen Behandlung von Singularitäten in elliptischen partiellen Differentialgleichungen. Bericht I-8810, Institut für Informatik, TU München, August 1988.

Ruede:1989:LCE

[Rüd89] U. Rüde. Local corrections for eliminating the pollution effect of reentrant corners. In J. Mandel, editor, Proceedings of the Fourth Copper Mountain Conference on Multigrid Methods, April 9-14, 1989, pages 365–382. SIAM, Philadelphia, PA, USA, 1989.

Ruede:1991:AHO

[Rüd91a] U. Rüde. Adaptive higher order multigrid methods. In W. Hackbusch and U. Trottenberg, editors, Proceedings of the Third European Conference on Multigrid Methods, October 1-4, 1990, pages 339–351. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, 1991. International Series of Numerical Mathematics, Vol. 98.

Ruede:1991:ERT

[Rüd91b] U. Rüde. Extrapolation and related techniques for solving elliptic equations. Bericht I-9135, Institut für Informatik, TU München, September 1991. URL

                                        file://www.tu-chemnitz.de/
                                        pub/Local/mathematik/Ruede/
                                        extra_rel.ps.Z.

Ruede:1992:DSM

[Rüd92a] U. Rüde. Data structures for multilevel adaptive methods and iterative solvers. Bericht I-9217, Institut für Informatik, TU München, May 1992. URL

                                        file://www.tu-chemnitz.de/
                                        pub/Local/mathematik/Ruede/
                                        data_structures.ps.Z.

Ruede:1992:HBE

[Rüd92b] U. Rüde. The hierarchical basis extrapolation method. SIAM J. Sci. Stat. Comput., 13(1):307– 318, January 1992. CODEN SI-JCD4. ISSN 0196-5204. Proceedings of the First Copper Mountain Conference on Iterative Methods, April 1-5, 1990, T. Manteuffel ed.

Ruede:1992:REF

[Rüd92c] U. Rüde. On the robustness and efficiency of the fully adaptive multigrid method. In A. Quarteroni, editor, Proceedings of the Sixth International Conference on Domain Decomposition in Science and Engineering, Como, Italy, June 15-19, 1992, pages ??–?? (of xxii + 484). Amer. Math. Soc., Providence, RI, USA, 1992. ISBN 0-8218-5158-6. LCCN QA402.2 .I55 1992.

Ruede:1992:VCF

[Rüd92d] U. Rüde. On the V-cycle of the fully adaptive multigrid method. Bericht I-9215, Institut für Informatik, TU München, May 1992. URL

file://www.tu-chemnitz.de/
pub/Local/mathematik/Ruede/

kiel93.ps.Z. to be published in the proceedings of the 9th GAMM Seminar, Kiel, January 22–24, 1993.

Ruede:1993:DAT

[Rüd93a] U. Rüde. Data abstraction techniques for multilevel algorithms. In Proceedings of the GAMM-Seminar on Multigrid Methods, Sept. 21 – 25, 1992 in Gosen, Germany. Institut für Angewandte Analysis und Stochastik, Berlin, Germany, 1993. URL

file://www.tu-chemnitz.de/
pub/Local/mathematik/Ruede/

gosen92.ps.Z. Report 5, ISSN 0942–9077.

Ruede:1993:EEB

[Rüd93b] U. Rüde. Error estimators based on stable splittings. Submitted to the proceedings of the 7th International Conference On domain Decomposition, Penn State University, 1993. URL

                                        file://www.tu-chemnitz.de/
                                        pub/Local/mathematik/Ruede/
                                        ddm7.ps.Z.

Ruede:1993:ETC

[Rüd93c] U. Rüde. Extrapolation techniques for constructing higher order finite element methods. Bericht I-9304, Institut für Informatik, TU München, 1993. URL

                                        file://www.tu-chemnitz.de/
                                        pub/Local/mathematik/Ruede/
                                        extrapolation.ps.Z.

Ruede:1993:MCT

[Rüd93d] U. Rüde. Mathematical and computational techniques for multilevel adaptive methods, volume 13 of Frontiers in Applied Mathematics. SIAM, Philadelphia, PA, USA, 1993. ISBN 0-89871-320-X. xii + 140 pp. LCCN QA377 .R87 1993.

Ruede:1993:MES

[Rüd93e] U. Rüde. Multilevel, extrapolation, and sparse grid methods. SFB Bericht 342/10/93 A/ I-9319, Institut für Informatik, TU München, July 1993. URL

                                        file://www.tu-chemnitz.de/
                                        pub/Local/mathematik/Ruede/

emg93.ps.Z. to appear in the Proceedings of the European Conference on Multigrid Methods, Amsterdam, July 6–9, P. Hemker and P. Wesseling eds.

Rude:1993:FAM

[Rüd93f] Ulrich Rüde. Fully adaptive multigrid methods. SIAM Journal on Numerical Analysis, 30(1): 230–248, February 1993. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Ruede:1994:MAI

[Rüd94] U. Rüde. On the multilevel adaptive iterative method. SIAM J. Sci. Stat. Comput., 15, 1994. CO-DEN SIJCD4. ISSN 0196-5204. also available as TU-Bericht I-9216, and published in the Preliminary Proceedings of the 2nd Copper Mountain Conference on Iterative Methods, April 9–14, 1992, ed. T. Manteuffel, University of Colorado at Denver.

Ruede:2002:CEP

[Rue02] U. Ruede. Computational engineering programs at the University of Erlangen-Nuremberg. Lecture Notes in Computer Science, 2331:852–??, 2002. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer-
ny.com/link/service/series/
0558/bibs/2331/23310852.htm;
 http://link.springer-ny.com/
link/service/series/0558/papers/
2331/23310852.pdf.

Ruede:1986:TSM

[RZ86a] U. Rüde and C. Zenger. On the treatment of singularities in the multigrid method. In W. Hackbusch and U. Trottenberg, editors, Lecture Notes in Mathematics 1228: Multigrid Methods II, Proceedings of the Conference Held at Cologne, October 1-4, 1985, pages ??– ?? (of vi + 335). Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1986. ISBN 0-387-16491-X. LCCN QA3 .L35 v.1228. DM50.00.

Ruede:1986:WMM

[RZ86b] U. Rüde and C. Zenger. A workbench for multigrid methods. Bericht I-8607, Institut für Informatik, TU München, May 1986.

Ruede:1992:TSF

[RZ92] U. Rüde and C. Zenger. On the treatment of singularities in the finite element method. Bericht I-9220, Institut für Informatik, TU München, August 1992.

Slavkovsky:1990:SBK

[SR90] P. Slavkovsky and U. Rüde. Schnellere Berechnung klassischer Matrix-Multiplikationen. SFB Bericht 342/17/90, Institut für Informatik, TU München, September 1990.

Zenger:1988:BSE

[ZFR88] C. Zenger, R. Fößmeier, and U. Rüde. Betriebssystem- und Software-Engineering-Aspekte bei parallelen Algorithmen. Kerntechnik, 52(2):120–125, 1988. CO-DEN KERNEU. ISSN 0004-7198, 0932-3902, 0368-5276.